given that the density of a solution of 64.5% C3H8O is 0.8752g/ml, what volume of this solution would contain 1.558x10^23 C3H8O molecules?
Answer #1

use:

number of mol = number of molecules / Avogadro’s number

number of mol = 1.558*10^23 / 6.022*10^23 mol

number of mol = 0.2587 mol

Molar mass of C3H8O,

MM = 3*MM(C) + 8*MM(H) + 1*MM(O)

= 3*12.01 + 8*1.008 + 1*16.0

= 60.094 g/mol

use:

mass of C3H8O,

m = number of mol * molar mass

= 0.2587 mol * 60.09 g/mol

= 15.55 g

This is 64.5 % of mass of solution

Use:

mass % = mass of C3H8O * 100 / mass of solution

64.5 = 15.55 * 100 / mass of solution

mass of solution = 24.11 g

Now use:

density = mass of solution / volume of solution

0.8752 g/mL = 24.11 g / volume of solution

volume of solution = 27.5 mL

Answer: 27.5 mL